OpenStudy (anonymous):

Solve (x+16)^2 + 18(x+18) = (x + 16)(x+18)

OpenStudy (ipwnbunnies):

Expand everything out, combine like terms. Get the x terms on one side and the constants on the other. Solve for x.

OpenStudy (anonymous):

Imma type the first part then I will continue solving. x^2+256+18x+324=x^2+18x+16x+288

OpenStudy (anonymous):

here is the next part x^2+18x+580=x^2+34x+288

OpenStudy (anonymous):

next x^2+18x+292=x^2+34x

OpenStudy (anonymous):

then x^2-16x+292=x^2

OpenStudy (anonymous):

(x+16)^2 + 18*(x+18)-(x + 16)*(x+18)=0 expanding--> 16*x+292 = 0 x = -73/4

OpenStudy (anonymous):

yeah that's the other way of doing it. I guess I was taught different lol

OpenStudy (anonymous):

the rest of mine was -16x+292=0 -16x=292 x= -18.25 or -73/4

OpenStudy (anonymous):

You are right pandaluver44, I just wanted to share my solution...